Find the volume of the solid of revolution by rotating the region formed by y=3-x2 and y=2 about the line y=2


1
Expert's answer
2021-07-26T16:07:12-0400

The given region in cartesian plan ooks like -



If this is rotated about y=2 , we get -


we have -




For ease of calculation, it is convenient if we shift this solid down so the axis of rotation falls along the X-axis:





Note that the radius (relative to the X-axis) of this shifted volume is

equal to "x^{2}-1" for "x\in[-1,+1]"


and we can slice this solids into thin disk, each with a thickness of "\delta" , so that each disc has volume of -


"{\delta}prr^{2}=\pi{\delta}(x^{2}-1)^{2}"




With very small values of "{\delta}"

 the sum of the volumes of all such disks will be the volume of the rotated solid.


We can evaluate this sum with "{\delta}" "\to" 0


 using the integral:


"=\int_{-1}^{1}{\pi}(x^{2}-1)^{2}dx"


"={\pi}\int_{-1}^{1}x^{4}-2x^{2}+1dx"


"={\pi}(\dfrac{x^5}{5}" "-2\dfrac{x^{3}}{3}" "+x)|_{-1}^{1}"


"=" "{\pi}(\dfrac{3-10+15}{15}-\dfrac{({-3})+10-15}{15})=\dfrac{16}{15}\pi"